OpenStudy (pgreene008):

Express the complex number in trigonometric form. -4

OpenStudy (pgreene008):

will fan and medal

myininaya (myininaya):

so first look on the unit circle where do you have there is a real part but no imaginary part

myininaya (myininaya):

what point

myininaya (myininaya):

you could actually two points from the unit circle but we only need one

OpenStudy (pgreene008):

ok

OpenStudy (pgreene008):

so what do i do?

myininaya (myininaya):

what point on the unit circle has a real part but no imaginary part

OpenStudy (pgreene008):

i guess all of them are real with no imaginaries, right?

myininaya (myininaya):

no

myininaya (myininaya):

the x-axis is the real number axis and the y-axis is the imaginary axis in this scenario

myininaya (myininaya):

(1,0) and (-1,0) both have a real but no imaginary part also known as 1+0i or -1+0i

myininaya (myininaya):

your number given is just a multiply of that

OpenStudy (pgreene008):

oh ok

myininaya (myininaya):

what angle has cosine is 1 and sine is 0?

OpenStudy (pgreene008):

90 degrees

myininaya (myininaya):

at 90 degrees we have cosine is 0 and sine is 1

OpenStudy (pgreene008):

oh wait, yeah youre right. its 0 degrees

myininaya (myininaya):

cool stuff... \[1+0i=\cos(0)+\sin(0)i\] and multiply both sides by -4 or... you could have gone with \[-1+0i=\cos(\pi)+\sin(\pi)i\] and instead multiply both sides by 4

OpenStudy (pgreene008):

so would it be: 4(cos 0° + i sin 0°) ?

myininaya (myininaya):

that is one possible answer out of many

myininaya (myininaya):

but yes that works

myininaya (myininaya):

i'm sorry

myininaya (myininaya):

I missed you didn't put the -4 in

myininaya (myininaya):

-4(cos(0)+i sin(0))

myininaya (myininaya):

two answers out of many -4(cos(0)+isin(0)) 4(cos(pi)+isin(pi))

myininaya (myininaya):

it also depends on what they want do they want r>0 the magnitude number?

myininaya (myininaya):

did they give any restrictions on r or the angle we choose

OpenStudy (pgreene008):

these were my possible answers: 4(cos 0° + i sin 0°) 4(cos 180° + i sin 180°) 4(cos 90° + i sin 90°) 4(cos 270° + i sin 270°)

myininaya (myininaya):

oh well we definitely have your choice above instead of using pi you could use...

myininaya (myininaya):

180 deg...

myininaya (myininaya):

pi rad=180 deg

myininaya (myininaya):

\[-1+0i=\cos(180^ \circ)+\sin(180^\circ) i\] multiply both sides by 4